Suppose f(x)=4x-4.
f(-2)=
f(x)+f(-2)=

Answer by Edwin McCravy(20055)


It's just a matter of plugging in what's in the parentheses of f( ) for x.

Suppose f(x)=4x-4.

f(-2)=

You just plug in -2 for x in

f(x)=4x-4.

f(-2)=4(-2)-4 = -8-4 = -12

-------------------

f(x)+f(-2)= 

You just plug in (4x-4) for f(x) and (-12) for f(-2)

f(x)+f(-2)= (4x-4)+(-12) = 4x-4-12 = 4x-16

Edwin